OpenStudy (anonymous):

Find the value of k so that the points are /sqrt {274} units apart. (-6, -8) (k, 7)

OpenStudy (anonymous):

\[ 274= (k+6)^2 + (7+8)^2 \] Solve for k

OpenStudy (anonymous):

After expansion and simplification, the equation above becomes \[ k^2+12 k-13=0\\ (k-1) (k+13)=0\\ k=1\\ k=-13 \]